OpenStudy (anonymous):

If f(x) = 5x – 4 and g(x) = 6x + 3, what is f(g(x))?

OpenStudy (anonymous):

f(x)=5x-4 f(g(x))=5g(x)-4 f(6x+3)=5(6x+3)-4 f(g(x))=30x+15-4=30x+11

OpenStudy (anonymous):

thanks